My question is
Without graphing identify the slope and the Y intercept
-3x + 5y = 5
please help with a simple understandable answer

Answer by Alan3354(69443)

Without graphing identify the slope and the Y intercept
-3x + 5y = 5
-------------------
You'll see this often.
To find the slope, m, and the y-intercept, put the equation in slope-intercept form: y = mx + b
That just means solve for y.
-3x + 5y = 5
5y = 3x + 5
y = (3/5)x + 1
The slope is the coefficient of x: m = 3/5
The y-intercept is +1, or the point (0,1)